# Accelerating Tactile Internet with QUIC: An Exploration of its Security and Privacy Attacks

### v1

@@ IV Potential Mitigation Strategies
This motivated the IETF to explore and encrypt the handshake messages apart from just the application data. In a work-in-progress document [^1] https://datatracker.ietf.org/doc/draft-ietf-tls-esni/16/, IETF proposes the inclusion of the Encrypted Client Hello (ECH) extension into the TLS to mitigate the security risks and improve the privacy of the Internet as a whole. The goal of ECH is to protect all sensitive handshake parameters such as preshared key, server name, etc. by encrypting the entire ClientHello, thereby closing the gap left by TLS 1.3 and Encrypted Server Name Identification (ESNI), the predecessor of ECH.
An extensive study was performed by Bhargavan et al. [13] to prove that TLS 1.3 with ECH indeed ensures basic security goals (like confidentiality, authenticity etc.), handshake privacy, and downgrade resistance. They provide a formal analysis of the privacy properties of TLS 1.3 in the presence of ECH using a symbolic protocol analyzer, named “ProVerif” to justify the security of ECH. However, the authors didn’t study privacy-preservation of ECH in the presence of traffic analysis. Therefore, any attack that relies on observing traffic patterns including the length of application messages and their timings would be something interesting to study in the future.
In another recent study [14], Trevisan et al. explore whether an attacker can recover the domain names of websites visited by a user in the presence of ECH by passively observing the traffic traces. The study reveals that simple machine learning models are capable of recovering a user’s visited websites with high accuracy. Their experiments reveal that ECH is only able to protect $\approx 7.5\%$ flows. In the end, they also evaluate the efficacy of standard padding-based techniques to thwart attacks on ECH. However, it is observed that the attackers are still able to gain a sufficiently large amount of information with just a minimal decrease in the effectiveness of the attacks launched. Hence, as a future work, it is important to improve ECH to increase its resistance to traffic analysis. Designing more robust techniques is necessary for offering better protection to end-users in terms of privacy.
While ECH can be a potential mitigation strategy against some of the security attacks (discussed in see: § II), Oblivious HTTP [^2] https://datatracker.ietf.org/doc/draft-ietf-ohai-ohttp/ and MASQUE (Multiplexed Application Substrate over QUIC Encryption) [^3] https://datatracker.ietf.org/doc/draft-ietf-masque-connect-ip/ are other potential work in progress technologies for the privacy attacks highlighted in § III. Private relay [15] offers privacy-preserving communication by establishing multiple proxy servers between the client and the server. This ensures that no attacker has visibility over both the user’s IP address and their access history. In this way, it separates the client’s IPs from servers and uses multi-hop MASQUE for proxying IP packets. On the contrary, Oblivious HTTP offers to send encrypted HTTP messages from a client to a gateway via a trusted relay service. This ensures that nobody sees the client’s IP address as well as the plaintext HTTP message remains hidden from the relay resources.
Overall, as a potential future work, researchers should consider designing privacy-preserving extensions or modifications to the QUIC protocol in order to mitigate potential privacy risks, such as by employing encryption for sensitive metadata and exploring differential privacy techniques. It is also crucial to integrate multi-factor authentication to fortify the trustworthiness of QUIC connections.
V Conclusion
The Tactile Internet is expected to bring a paradigm shift in all spheres of the connected society. However, this is only possible with tight integration of the network infrastructure, transport, and application (Tactile Internet) layers. Moreover, addressing the security and privacy challenges is crucial for the success of the Tactile Internet. In this article, we envisioned a QUIC-accelerated Tactile Internet that leverages 6G as the underlying network infrastructure for achieving low-latency, resilient global connectivity. We explored this ecosystem from a security and privacy perspective and presented several attacks that can make Tactile Internet vulnerable by targeting QUIC. To this end, we presented a few potential mitigation strategies that can be the future directions to plug the present security and privacy gap in this ecosystem, to help the Tactile Internet transition the connected society into a more sustainable future.

### v3

@@ 4.1. Encrypting Handshake Metadata
Despite TLS1.3 being encrypted, certain information are exchanged in plaintext during the initial connection establishment phase, allowing the eavesdroppers to learn various crucial metadata and other sensitive information like Server Name Identification (SNI) extension (see: § 2) available in the CHLO message. This drawback particularly motivated the IETF to encrypt the handshake messages apart from just the application data. IETF proposes the inclusion of the Encrypted Client Hello (ECH) extension (Rescorla et al., 2025) into the TLS to mitigate the security risks and improve the privacy of the Internet as a whole. The goal of ECH is to protect all sensitive handshake parameters such as preshared key, server name, etc. by encrypting the entire CHLO, thereby closing the gap left by TLS 1.3 and Encrypted Server Name Identification (ESNI), the predecessor of ECH. A recent study (Bhargavan et al., 2022), shows that ECH, indeed ensures basic security goals (like confidentiality, authenticity etc.), while offering stronger downgrade resistance handshake privacy and broader metadata protection. Despite these benefits offered by ECH, several practical challenges still persist:
  • Partial Deployment and Downgrade Risks: Many servers still do not support ECH, leading to fallback to unencrypted handshakes. Attackers can exploit this partial deployment by forcing fallback, exposing sensitive metadata.
  • Middlebox Interference: Numerous middleboxes incorrectly assume visibility into TLS extensions. RFC 9325 (Sheffer et al., 2022) outlines middlebox tolerance recommendations, but real-world compliance remains limited.
  • Traffic Analysis on Encrypted Handshakes: A recent study (Trevisan et al., 2023) reveals that simple machine learning models are capable of recovering a the domain names of a user’s visited websites even in the presence of ECH with very high accuracy. Attackers can infer the target server or user activity through traffic features such as handshake size, packet timing, and record counts, even when ECH is employed. Even in the presence of standard padding-based techniques, attackers are still able to gain a sufficiently large amount of information
  • Dependency on DNS Security: ECH relies on the correct and secure delivery of public keys via DNS SVCB/HTTPS (Schwartz et al., 2023) records. Attacks on DNS privacy, such as via DNS cache poisoning or compromised resolvers, undermine ECH’s guarantees.
Future research should prioritize developing efficient, traffic-aware padding and shaping mechanisms to better protect handshake messages against traffic analysis attacks. In parallel, it is crucial to enforce integrated defenses that span both DNS resolution and TLS handshakes, ensuring that metadata privacy is preserved throughout the resolution and connection establishment phases; specifically, tight coupling between DNS-over-QUIC (DoQ) (Huitema et al., 2022) and TLS encryption is necessary to prevent potential leakage through resolution paths. Additionally, improving the deployability of ECH remains a pressing concern, motivating the exploration of progressive deployment strategies such as enabling clients to attempt encrypting CHLO messages even in the absence of explicit server-side ECH advertisements.
4.2. Privacy-preserving Proxying
The IETF has developed complementary protocols to strengthen user privacy beyond the transport layer. Oblivious HTTP (OHTTP) (RFC 9484 (Thomson and Wood, 2024)) decouples client IP addresses from request payloads by relaying encrypted HTTP requests through trusted intermediaries. This design ensures that no single entity can observe both the sender’s identity and the content of the request. Similarly, MASQUE (Multiplexed Application Substrate over QUIC Encryption) (Schinazi, 2025) enables tunneling of arbitrary IP and UDP flows over QUIC connections, thereby supporting privacy-preserving proxying, virtual private network (VPN) services, and flexible secure tunneling for broader Internet applications.
Despite these advances, OHTTP and MASQUE introduce several operational and security challenges. First, OHTTP assumes that relays and target gateways do not collude; however, in adversarial environments, relay compromise or collusion can severely undermine user privacy guarantees. Second, both OHTTP and MASQUE proxies may be exploited for amplification attacks unless they implement robust rate-limiting, authentication, and authorization mechanisms to control resource usage. Third, additional relaying inevitably introduces latency and bandwidth overheads, raising trade-offs between enhanced privacy and acceptable performance metrics, particularly for real-time or mobile applications.
4.3. Resilience Against Resource Exhaustion
While mechanisms such as ECH and privacy-preserving proxying via OHTTP and MASQUE strengthen user privacy, they also increase server-side computational and memory demands during connection establishment. As outlined in § 2.6 and  2.7, this heightened state maintenance makes QUIC endpoints more vulnerable to resource exhaustion and DoS attacks, where adversaries exploit expensive handshake operations to overwhelm servers.
To mitigate these risks without negating the privacy benefits of ECH and OHTTP, several approaches can be adopted. Refining QUIC’s stateless RETRY mechanisms can enable earlier validation of client authenticity during handshake attempts, ensuring that malicious clients are filtered with minimal server resource commitment. Further, under extreme load conditions, servers may employ graceful privacy degradation strategies, such as offering opportunistic encryption modes that maintain basic confidentiality without the full overhead of ECH or multi-hop proxying, thus ensuring continued service availability without complete privacy forfeiture. Together, these approaches ensure that privacy enhancements do not inadvertently compromise the operational resilience of QUIC deployments.
5. QUIC in the Post-Quantum Era
While the mitigation strategies discussed in § 4 address current threats to QUIC security and privacy, an important future challenge is ensuring resilience against quantum-capable adversaries. Cryptographically relevant quantum computers could compromise classical public key systems, weakening handshake security and long-term confidentiality. To proactively mitigate this risk, the IETF is standardizing hybrid key exchange approaches that combine classical and post-quantum algorithms (Stebila et al., 2025), while also developing broader deployment guidelines for transport and application-layer protocols (Banerjee et al., 2025; Reddy.K and Tschofenig, 2025).
Integrating post-quantum cryptography (PQC) into QUIC presents several operational concerns. Post-quantum key encapsulation mechanisms (KEMs) notably increase handshake sizes, exacerbating risks of resource exhaustion attacks (§ 2.62.7) by amplifying computational and memory burdens on servers. Additionally, hybrid handshakes may create new traffic analysis vectors (§ 3.2), allowing adversaries to distinguish handshake types based on packet sizes, flow timings, or retransmission patterns. Such metadata leakage could undermine the privacy goals of mechanisms like ECH, especially if hybrid and classical flows become easily identifiable.
Mitigating these risks requires compact hybrid handshake designs that minimize handshake expansion while preserving cryptographic strength. Furthermore, formal modeling of handshake privacy under hybrid and post-quantum settings is necessary to evaluate indistinguishability guarantees. Finally, traffic shaping and padding mechanisms must evolve to equalize observable characteristics across classical, hybrid, and future post-quantum-only sessions, preventing adversarial profiling. Ensuring that QUIC remains resilient in a post-quantum world will demand careful balancing of cryptographic security, metadata privacy, and operational scalability. Addressing these engineering challenges proactively is essential to maintaining QUIC’s role as a secure, future-proof transport protocol.
6. Conclusion
This paper revisited the evolving security and privacy challenges faced by QUIC, analyzing vulnerabilities in handshake metadata exposure, traffic analysis, and resilience under adversarial conditions. We critically evaluated emerging mitigation strategies, including encrypted handshake mechanisms, privacy-preserving proxy architectures, and resource exhaustion defenses, highlighting both the advances made and the challenges that persist in real-world deployments. Additionally, we examined the operational complexities introduced by post-quantum cryptographic handshakes, identifying new risks related to handshake amplification and metadata leakage. Our findings reveal that while significant progress has been made, important deployment gaps persist despite theoretical advancements. Bridging this gap will require continued research into adaptive privacy mechanisms, traffic pattern obfuscation, cross-layer resilience, and post-quantum readiness. Proactive engineering towards robust, privacy-preserving, and quantum-resilient QUIC is critical to securing the next-generation Internet.